PLEASE HELP ME SOLVE THIS EQUATION!!!
4(2x-3)-3(3x-10)=11
Thankyou =)

Answer by venugopalramana(3286)

4(2x-3)-3(3x-10)=11
8x-12-9x+30=11
-x=11+12-30=-7
x=-7/-1=7
x=7